# Chapter 0 Hadron Production Processes

### v1

@@ 9 Double-Pion Production on the Nucleon
In certain energy regions, the $\pi N\to 2\pi N$ reaction accounts for up to 50% of the $\pi N$ inelasticity, as seen from Fig. 30. Therefore, this production channel must be included in any CC-PWA approach, as e.g., practiced in GiM. An improved and considerably extended description of double-pion production within our coupled channels scheme was started recently, and first results are found in Ref. [184]. The inclusion of multi-meson configurations into a coupled channels approach is a highly non-trivial exercise in three-body dynamics. In view of the complexities, physically meaningful approximations are necessary, retaining the essential dynamical aspects but making numerical calculations feasible. For that goal, the ansatz used in Ref. [184] relies on an isobar description of intermediate two-pion configurations and their decay into the final double-pion states on the mass shell. The derived processes contributing to the T-matrix of double-pion production on the nucleon in that energy region are depicted in Fig. 30.
@@
This approach allows for the direct analysis of the $2\pi N$ experimental data. Since the corresponding Dalitz plots are found to be strongly non-uniform, it is natural to assume that the main effect to the reaction comes from the resonance decays into isobar sub-channels [185]. The most important contributions are expected to be from the intermediate $\sigma N$, $\pi\Delta$, and $\rho N$ states. The analysis of the $\pi N\to 2\pi N$ reaction would therefore provide very important information about the resonance decay modes into different isobar final states. The much richer baryon spectrum found in LQCD simulations [186, 187], the functional DSE approaches [188], and the CQM results [189, 190] that are observed in scattering experiments indicate the necessity for broader investigations, including a larger class of reaction channels. Experimentally, most of the non-strange baryonic states have been identified from the analysis of the elastic $\pi N$ data [181, 98]. As pointed out in [189] the signal of excited states with a small $\pi N$ coupling will be suppressed in the elastic $\pi N$ scattering. As a solution to this problem, a series of photoproduction experiments has been done to accumulate enough data for the study of the nucleon excitation spectra. However, the results from the photoproduction reactions are still controversial. While recent investigations of the photoproduction reactions presented by the BnGa group [191] reported indications for some new resonances, not all of these states are found in other calculations [182]. This raises a question about independent confirmation for the existence of such states from the investigations of other reactions.
@@
Because of the smallness of the electromagnetic couplings, the largest contribution to the resonance self-energy comes from the hadronic decays. If the $N^{\ast}\to\pi N$ transition is small, one can expect sizable resonance contribution into remaining hadronic decay channels. As a result, the effect from the resonance with a small $\pi N$ coupling could still be significant in the inelastic pion-nucleon scattering: here the smallness of resonance coupling to the initial $\pi N$ states could be compensated by the potentially large decay branching ratio to other different inelastic final states. Such a scenario is realized, e.g., in the case of the well-known $N^{\ast}(1535)$ state. While the effect from this resonance to the elastic $\pi N$ scattering is only moderate at the level of the total cross section, its contribution to the $\pi N\to\eta N$ channel turns out to be dominant [180]. Since the $\pi N\to 2\pi N$ reaction could account for up to 50% of the total $\pi N$ inelasticity, this channel becomes very important not only for the investigation of the properties of already known resonances but also for the search for the signals of possibly unresolved states.
@@
Figure 30: The processes contributing to double-pion production T-matrix are depicted diagrammatically: (a) and (b) production through the $\sigma$-isobar, (c) and (d) production through the $\Delta^{0}$-isobar. Symmetrization is indicated.
@@
Another important issue in studies of the $2\pi N$ channel is related to the possibility to investigate cascade transitions like ${N^{\ast}}^{\prime}\to\pi N^{\ast}\to\pi\pi N$, where a massive state ${N^{\ast}}^{\prime}$ decays via intermediate excited $N^{\ast}$ or $\Delta^{\ast}$. It is interesting to check whether such decay modes are responsible for the large decay width of higher lying mass states. So far only the $\pi N^{\ast}(1440)$ isobar channel has been considered in a partial wave analysis (PWA) of the $\pi N\to 2\pi N$ experimental data [185].
@@
There are several complications in the coupled-channel analysis of $2\to 3$ transitions. The first one is the difficulty to perform the partial-wave decomposition of the three-particle state. The second complication is related to the issue of three-body unitarity. For a full dynamical treatment of the $2\to 3$ reaction, the Faddeev equations (named after Ludwig Dmitrievich Faddeev (1934-2017)) have to be solved. Although appropriate theoretical and numerical methods are known, the effort inhibits practical implementations. Here, a coupled-channel approach for solving the $\pi N\to 2\pi N$ scattering problem in the isobar approximation is used, as widely practiced. In this formulation, the $(\pi/\pi\pi)N\to(\pi/\pi\pi)N$ coupled-channel equations are reduced to effective two-body scattering equations, taking advantage of intermediate two-body isobar production. Such a description accounts by construction for the full spectroscopic strength of intermediate channels and, in addition, provides a considerable numerical simplification.
@@
Three-body unitarity leads to a relation between the imaginary part of the elastic scattering amplitude and the sum of the total elastic and inelastic cross sections by the well-known optical theorem. Since in the isobar approximation the pions in the $\pi\pi N$ channel are produced from the isobar sub-channels, all contributions to the total $\pi N\to\pi\pi N$ cross section are driven by the isobar production. The optical theorem can be fulfilled if all discontinuities in isobar sub-channels are taken into account. In the present work, the three-body unitarity is maintained up to interference terms between the isobar sub-channels.
@@
The first resonance energy region is of particular interest because of the sizable effect from $N^{\ast}(1440)$. The dynamics of the Roper resonance turns out to be rich because of the two-pole structure reported in earlier studies [192, 193], (see [181, 194, 195] for the recent status of the problem). The origin of the Roper resonance is also controversial. For example, the calculations in the Jülich model explain this state as a dynamically generated pole due to the strong attraction in the $\sigma N$ sub-channel. At the same time, the Crystal Ball Collaboration finds no evidence of strong $t$-channel sigma-meson production in their $\pi^{0}\pi^{0}$ data [196]. From the further analysis of the $\pi^{0}\pi^{0}$ production, the effect of the sigma meson was found to be small [197]. On the other hand, the $pp\to pp\pi^{0}\pi^{0}$ scattering experiment by the CELSIUS-WASA collaboration [198] finds the $\sigma N$ decay mode of the Roper resonance to be dominant.
@@
Figure 31: $\pi^{0}\pi^{0}$ differential cross sections for the reaction $\pi^{-}p\to\pi^{0}\pi^{0}n$ at fixed $t_{\pi^{0}\pi^{0}}=\cos{\theta_{\pi^{0}\pi^{0}}}$, shown in the upper left corner of the panels. Energy distributions for $-0.95\leq t_{\pi^{0}\pi^{0}}\leq+0.95$ are shown. The experimental data are from [197] The figure is adapted from Ref. [184].
@@
In the region of the Roper resonance, our calculations are able to describe the mass distributions rather satisfactorily. Also, in this region, the production strength is shifted to higher invariant masses $m_{\pi^{0}\pi^{0}}^{2}$. At the same time, a peak at small $m_{\pi^{0}\pi^{0}}^{2}$ becomes also visible. In the present calculations, the fit tends to decrease the magnitude of the $\pi\Delta$ production and compensate it by enhancing the strength into $\sigma N$. The obtained decay branching ratio of $N^{\ast}(1440)$ for the $\sigma N$ channel is about twice as large as for the $\pi\Delta$.
@@
Both the small peak at small and the broad structure at large invariant masses are well reproduced indicating an important interplay between the $\sigma N$ and $\pi\Delta$ production mechanisms. It is interesting that the isoscalar correlations in the $\pi\pi$ rescattering are also found to be necessary in order to reproduce the asymmetric shape of the mass distributions. Though the $\pi\Delta$ production gives rise to a two-peak structure only the first one at small $m_{\pi^{0}\pi^{0}}^{2}$ is visible at energies $1.4-1.468~{}\mathrm{GeV}$. Within the present calculation the second peak at high $m_{\pi^{0}\pi^{0}}^{2}$ is not seen because of the large $\sigma N$ contributions. In the present study $\pi^{0}\pi^{0}n$ production is calculated as a coherent sum of isobar contributions. Though the interference effects are important they are found to be very small at the level of the total cross sections.
@@
To simplify the analysis, the $S_{11}$ and $P_{11}$ $\pi N$ partial waves are directly constrained by the single energy solutions (SES) derived by GWU(SAID) [181]. The experimental data on the $\pi^{-}p\to\pi^{0}\pi^{0}n$ reaction are taken from [197]. These measurements provide high-statistics data on the angular distributions ${\rm d}\sigma/{d\Omega_{\pi\pi}}$ where $\Omega_{\pi\pi}$ is the scattering angle of the $\pi\pi$ pair (or the final nucleon in c.m.).
@@
The calculated $\pi^{0}\pi^{0}$ differential cross sections are shown in Fig. 31 and compared to the Crystal Ball data as a function of the c.m. energy. The measurements demonstrate a rapid rise of the cross sections at the energies $1.3-1.46~{}\mathrm{GeV}$, indicating the strong contribution coming from the Roper resonance, as also found in the GWU(SAID) [181] analysis.
@@
The invariant $\pi^{0}\pi^{0}$ mass distributions play a crucial role in the separation of the isobar contributions. The $\pi^{-}p\to\pi^{0}\pi^{0}n$ reaction close to threshold is dominated by the $\sigma N$ production due to the $t-$ channel pion exchange. The nucleon Born term contribution to the $\pi\Delta$ channel is found to be less significant. The decay branching ratios of $N^{\ast}(1440)$ are obtained as $R_{\sigma N}^{N(1440)}=27^{+4}_{-9}\,\%$ and $R_{\pi\Delta}^{N(1440)}=12^{+5}_{-3}\,\%$.
@@
The parameters extracted independently in the different approach of the BnGa group are $R_{\sigma N}^{N(1440)}=17^{+7}_{-7}\%$ and $R_{\pi\Delta}^{N(1440)}=21^{+8}_{-8}\%$, demonstrating that in spite of the visible difference in the central values, these quantities still coincide within error bars.
The double-pion channels are of high interest for $\chi$EFT. In [199], the Jülich group has been studying pion production off nucleons in the heavy baryon chiral perturbation theory to third order in the chiral expansion aiming at the determination of the low–energy constants. Most of the at that time available differential cross sections and angular correlation functions at low pion incident energies could be described together with total cross sections at higher energies. The contributions from the one-loop graphs were found to be essentially negligible once the dominant terms at second and third order are related to pion–nucleon scattering graphs with one pion were added. An interesting aspect is that the $\pi\pi N$ channels provide the possibility of extracting the pion–pion $S$–wave scattering lengths which otherwise are hard to access.
10 Double Pseudoscalar Meson Production Induced by Pions
1 Double-Pion Production by Pion Beams
Early studies of the dynamics of the $N\pi\pi$ final state were based on bubble chamber data on the reaction $\pi p\to\pi\pi N$ collected in non-polarized experiments by Berkeley, Saclay, and Rutherford laboratories (Fig. 32). 241,214 Bubble Chamber events have been analyzed in Isobar-model PWA at $W=1320-1930~{}\mathrm{MeV}$ by the Virginia Tech group led by Richard Arndt [185]. A summary of the number of Bubble Chamber events is given in Fig. 32.
Partial-wave inelasticities for elastic pion-nucleon scattering were determined with the aid of experimental data on $\pi N\to\pi\pi N$ processes in the beam-momentum range $300~{}\mathrm{MeV/c}<\mathrm{P_{beam}}<500~{}\mathrm{MeV/c}$ [200].
However, an obstacle for a much-desired joint PWA analysis is inconsistencies which at present inhibit merging $\pi N\to\pi N$ and $\pi N\to\pi\pi N$ databases. Problems exist specifically for small inelasticities (Fig. 33(left)), while the largest inelastic cross section related to the $P_{11}$ Roper resonance is in excellent agreement, being well described in SAID-SP06 for $\pi N\to\pi N$ [181] (Fig. 33(right)).
A complete analysis of $\gamma N\to\pi\pi N$ ideally would require fitting all data obtained with both pion and photon beams. Already a long while ago, Pontecorvo considered a joint analysis of pion-nucleon scattering and single pion production in nucleon-nucleon and pion-nucleon interactions [201] but a full scale analysis of that kind is still pending.

### v2

@@ 2 Double-Pion Production in Pion-Proton Reactions
In certain energy regions, the $\pi N\to 2\pi N$ reaction accounts for up to 50% of the $\pi N$ inelasticity. Therefore, this production channel must be included in any CC-PWA approach, as e.g., practiced in GiM Ref. [212, 83]. The Lagrangians of the interactions considered in double-$\pi^{0}$ production are displayed in Fig. 34. In view of the foregoing complexities, physically meaningful approximations are necessary. The guiding principle is to retain the essential dynamical aspects while making numerical calculations feasible. This goal is reached by an isobar description of intermediate two-pion configurations and their decay into the final double-pion states on the mass shell. The processes contributing to the T-matrix of $\pi^{0}\pi^{0}$ pair production on the nucleon in that energy region are depicted in Fig. 35.
Figure 34: The GiM Lagrangians underlying two-pion production on the nucleon. In the left column, the Lagrangian for $\pi\pi\leftrightarrow\sigma$, $\pi N\Delta$, and $\pi N^{*}\Delta$ interactions are shown, where in the bottom row the upper and lower choice of operator relates to $N^{*}$ of $J^{\pi}=\frac{1}{2}^{\pm}$. In the right column, the $\pi NN$ Lagrangian and the Lagrangians of mesons $\varphi\neq\pi$ interacting with a nucleon and $J^{\pi}=\frac{1}{2}^{\pm}$ resonances are displayed. The upper and lower choice of operator relates to mesons of negative and positive intrinsic parity, respectively.
@@
This approach allows for the direct analysis of the $2\pi N$ experimental data. Since the corresponding Dalitz plots have a strongly non-uniform structure, it is natural to assume that the main effect of the reaction comes from the resonance decays into isobar sub-channels [203]. The most important contributions are expected to be from the intermediate $\sigma N$, $\pi\Delta$, and $\rho N$ states. The analysis of the $\pi N\to 2\pi N$ reaction would, therefore, provide very important information about the resonance decay modes into different isobar final states. The much richer baryon spectrum found in LQCD simulations [213, 214], the functional DSE approaches [215], and the CQM results [216, 217] that are observed in scattering experiments indicate the necessity for broader investigations, including a larger class of reaction channels. Experimentally, most of the non-strange baryonic states have been identified from the analysis of the elastic $\pi N$ data [166, 111]. As pointed out in [216], the signal of excited states with a small $\pi N$ coupling will be suppressed in the elastic $\pi N$ scattering. As a solution to this problem, a series of photoproduction experiments has been conducted to accumulate enough data for the study of the nucleon excitation spectra. However, the results from the photoproduction reactions are still controversial. While recent investigations of the photoproduction reactions presented by the BnGa group [218] reported indications of some new resonances, not all of these states are found in other calculations [167]. This raises a question about the independent confirmation of such states from the investigations of other reactions.
@@
Because of the smallness of the electromagnetic couplings, the largest contribution to the resonance self-energy comes from the hadronic decays. If the $N^{\ast}\to\pi N$ transition is small, one can expect a sizable resonance contribution to the remaining hadronic decay channels. As a result, the effect of the resonance with a small $\pi N$ coupling could still be significant in inelastic pion-nucleon scattering. here, the smallness of the resonance coupling to the initial $\pi N$ states could be compensated for by the potentially large decay branching ratio to other inelastic final states. Such a scenario is realized, e.g., in the case of the well-known $N^{\ast}(1535)$ state. While the effect of this resonance on the elastic $\pi N$ scattering is only moderate at the level of the total cross section, its contribution to the $\pi N\to\eta N$ channel turns out to be dominant [165]. Since the $\pi N\to 2\pi N$ reaction could account for up to 50% of the total $\pi N$ inelasticity, this channel becomes very important not only for the investigation of the properties of already known resonances, but also for the search for the signals of possibly unresolved states.
@@
Figure 35: The processes contributing to double-pion production T-matrix are depicted diagrammatically: (a) and (b) production through the $\sigma$-isobar, (c) and (d) production through the $\Delta^{0}$-isobar. Symmetrization of the exit channels is indicated in, see (b) and (d).
@@
Another important issue in studies of the $2\pi N$ channel is the possibility of investigating cascade transitions like ${N^{\ast}}^{\prime}\to\pi N^{\ast}\to\pi\pi N$, where a massive state $N^{\ast}$$^{\prime}$ decays via intermediate excited $N^{\ast}$ or $\Delta^{\ast}$. The Lagrangians shown in Fig. 34 contain the proper interactions. Experimentally, such processes are clearly observed, as seen in Fig. 36 where Dalitz plots are shown for two-pion photo-production through the excitation of the $P_{33}(1232)$ Delta-resonance. It is interesting to check whether such decay modes are responsible for the large decay width of higher-lying mass states. So far, only the $\pi N^{\ast}(1440)$ isobar channel has been considered in a partial wave analysis (PWA) of the $\pi N\to 2\pi N$ experimental data [203].
Figure 36: Double-pion production through excitation of the $P_{33}(1232)$ Delta-resonance. The Dalitz plots were generated from CBELSA (left) and CLAS (right) photo-production data .
@@
The coupled-channel analysis of $2\to 3$ transitions is confronted with several complications. The first one is the difficulty of performing the partial-wave decomposition of the three-particle state. The second complication is related to the issue of three-body unitarity. For a full dynamical treatment of the $2\to 3$ reaction, the Faddeev equations must be solved. Although appropriate theoretical and numerical methods are known, the effort inhibits practical implementation. Here, a coupled-channel approach for solving the $\pi N\to 2\pi N$ scattering problem in the isobar approximation is used, as is widely practiced. In this formulation, the $(\pi/\pi\pi)N\to(\pi/\pi\pi)N$ coupled-channel equations are reduced to effective two-body scattering equations, taking advantage of intermediate two-body isobar production. Such a description accounts, by construction, for the full spectroscopic strength of intermediate channels and, in addition, provides a considerable numerical simplification.
Figure 37: Overview of elemental Feynman diagrams considered in the calculation of the 2-to-3-body amplitudes in double-pion production reactions. The incident photon may be replaced by an incoming hadron.
@@
Three-body unitarity leads to a relation between the imaginary part of the elastic scattering amplitude and the sum of the total elastic and inelastic cross sections, as described by the well-known optical theorem. Since, in the isobar approximation, the pions in the $\pi\pi N$ channel are produced from isobar sub-channels, all contributions to the total $\pi N\to\pi\pi N$ cross section are driven by isobar production. The optical theorem can be fulfilled if all discontinuities in isobar sub-channels are taken into account. In the present work, three-body unitarity is maintained up to interference terms between the isobar sub-channels. The calculation of the 2-to-3-body amplitude within this scheme is illustrated in Fig. 37.
@@
The first resonance energy region is of particular interest because of the sizable effect from $N^{\ast}(1440)$. The dynamics of the Roper resonance turn out to be rich because of the two-pole structure reported in earlier studies [219, 220], (see [166, 221, 222] for the recent status of the problem). The origin of the Roper resonance is also a matter of controversy. For example, the calculations in the Jülich model, explain this state as a dynamically generated pole due to the strong attraction in the $\sigma N$ sub-channel. At the same time, the Crystal Ball Collaboration finds no evidence of strong $t$-channel sigma-meson production in their $\pi^{0}\pi^{0}$ data [223]. From the further analysis of the $\pi^{0}\pi^{0}$ production, the effect of the sigma meson was found to be small [224]. On the other hand, the $pp\to pp\pi^{0}\pi^{0}$ scattering experiment by the CELSIUS-WASA collaboration [225] finds the $\sigma N$ decay mode of the Roper resonance to be dominant.
@@
Figure 38: $\pi^{0}\pi^{0}$ differential cross sections for the reaction $\pi^{-}p\to\pi^{0}\pi^{0}n$ at fixed $t_{\pi^{0}\pi^{0}}=\cos{\theta_{\pi^{0}\pi^{0}}}$, shown in the upper left corner of the panels. Energy distributions for $0.95\leq t_{\pi^{0}\pi^{0}}\leq+0.95$ are shown. The experimental data are from [224] The figure is adapted from Ref. [212].
@@
In the region of the Roper resonance, the calculations are able to describe the mass distributions quite satisfactorily. Also, in this region, the production strength shifts to higher invariant masses $m_{\pi^{0}\pi^{0}}^{2}$. At the same time, a peak at small $m_{\pi^{0}\pi^{0}}^{2}$ also becomes visible. In these calculations, the fit tends to decrease. The magnitude of the $\pi\Delta$ production compensates for it by enhancing the strength to $\sigma N$. The obtained decay branching ratio of $N^{\ast}(1440)$ for the $\sigma N$ channel is about twice as large as that for the $\pi\Delta$.
Figure 39: Two-pion production partial wave cross sections comparing GiM results to the Manley and GWU analyses as indicated. Adapted from Ref. [212].
@@
Both the small peak at small invariant masses and the broad structure at large invariant masses are well reproduced, indicating an important interplay between the $\sigma N$ and $\pi\Delta$ production mechanisms. Interestingly, the isoscalar correlations in the $\pi\pi$ rescattering are also found to be necessary in order to reproduce the asymmetric shape of the mass distributions. Though the $\pi\Delta$ production gives rise to a two-peak structure, only the first one at small $m_{\pi^{0}\pi^{0}}^{2}$ is visible at energies $1.4-1.468~\mathrm{GeV}$. The second peak at high $m_{\pi^{0}\pi^{0}}^{2}$ is not seen because of the large $\sigma N$ contributions. In the current approaches $\pi^{0}\pi^{0}n$ production is calculated as a coherent sum of isobar contributions. Though the interference effects are important, they are found to be very small at the level of the total cross sections.
@@
To simplify the analysis, the $S_{11}$ and $P_{11}$ $\pi N$ partial waves are directly constrained by the single energy solutions (SES) derived by GWU(SAID) [166]. The experimental data on the $\pi^{-}p\to\pi^{0}\pi^{0}n$ reaction are taken from [224]. These measurements provide high-statistics data on the angular distributions ${\rm d}\sigma/{d\Omega_{\pi\pi}}$, where $\Omega_{\pi\pi}$ is the scattering angle of the $\pi\pi$ pair (or the final nucleon in c.m.).
@@
The calculated $\pi^{0}\pi^{0}$ differential cross sections are shown in Fig. 38 and compared to the Crystal Ball data as a function of the c.m. energy. The measurements demonstrate a rapid rise in the cross sections at the energies $1.3-1.46~\mathrm{GeV}$, indicating a strong contribution from the Roper resonance, as also found in the GWU(SAID) [166] analysis.
@@
The invariant $\pi^{0}\pi^{0}$ mass distributions play a crucial role in the separation of the isobar contributions. The $\pi^{-}p\to\pi^{0}\pi^{0}n$ reaction close to the threshold is dominated by the $\sigma N$ production due to the $t$ channel pion exchange. The nucleon Born term contribution to the $\pi\Delta$ channel is found to be less significant. The decay branching ratios of $N^{\ast}(1440)$ are obtained as $R_{\sigma N}^{N(1440)}=27^{+4}_{-9}\,\%$ and $R_{\pi\Delta}^{N(1440)}=12^{+5}_{-3}\,\%$.
@@
The parameters extracted independently using the different approach of the BnGa group [226] are $R_{\sigma N}^{N(1440)}=17^{+7}_{-7}\%$ and $R_{\pi\Delta}^{N(1440)}=21^{+8}_{-8}\%$, demonstrating that, despite the visible differences in the central values, these quantities still coincide within the error bars.
In Fig.39, two-pion partial wave cross sections are compared for three different CC approaches. Although the GiM approach differs in detail considerably from the Manley and GWU-SAID descriptions, the results are in favorable overall agreement, thus confirming that the essential physics features are covered in the CC scheme.
Figure 40: Angular distributions of pions in the nucleon-pion (left) and nucleons in the pion-pion (right) helicity frames for the $\pi^{-}p\to h\pi^{+}\pi^{-}$ (upper) and $\pi^{-}p\to p\pi^{0}\pi^{-}$ (lower) reaction channels. The helicity frame is indicated by the subscript labels, while the angle between the given particles in that frame is denoted by superscript labels. The z-axis of the helicity frame is chosen opposite to the neutron (upper panel) and the proton (lower panel) directions. Color (gray) curves display various final-state contributions (indicated in the legend). The figure is adapted from Ref. [227].
Double-pion production has been part of the research program of the HADES@GSI experiment. The GSI pion beam facility was used to measure excitation functions and angular distributions for $\pi^{-}p\to h\pi^{+}\pi^{-}$ and $\pi^{-}p\to h\pi^{0}\pi^{-}$ reactions at four different pion beam momenta: 0.650, 0.685, 0.733, and $0.786~\mathrm{GeV/c}$ [227]. The primary goal was to study the very specific question of which of the resonances in the second resonance region around $1600~\mathrm{MeV}$ are dominant in $N\rho$ interactions. In Fig. 40, angular distributions in the $n\pi^{+}\pi$- and $p\pi^{0}\pi^{-}$ final channels are shown. The BnGa framework was used to study the partial wave and the isotopic channel composition of the measured angular distributions. Two major conclusions are drawn: First, the double-pion cross section is influenced by interference effects between isospin $I=1/2$ and $I=3/2$ amplitudes, which are constructive for $n\rho^{0}$ and destructive in the $p\rho^{-}$ case. Second, identify $N(1520)\frac{3}{2}^{-}$ with a branching ratio of $12,2\%$ as the dominant state for $N\rho$ self-energies.